A reservoir of glycerin has a mass of 1200 kg and a volume of 0.952 m3



. Find the



glycerin’s weight (W), density (𝜌), specific weight (𝛾), and specific gravity (SG).

1
Expert's answer
2021-11-18T14:59:46-0500

glycerin’s weight (W)

"F= W = ma = (1200)(9.81) = 11770N or 11.77kN"

density (𝜌)

"\rho = \frac{m}{V} = \frac{1200}{0.952} =1261kg/m^3"


specific weight (𝛾)

"\gamma = \frac{W}{V} =\frac{11.77}{0.952} =12.36kN/m^3"


specific gravity (SG)

"SG = V3,./yflzo Mac = \frac{12.36}{9.81} =1.26"